OpenStudy (anonymous):

Use estimation to determine which of the following quotients you expect to be less than 1. four-fifteenths divided by six−fifteenths seven-eighths divided by three-eighths nine-elevenths divided by one-eleventh five-sevenths divided by one-seventh

OpenStudy (anonymous):

@juliannescriv @math&ing001

OpenStudy (math&ing001):

(4/15) / (6/15) = 4/6 = 0.6666 < 1